Ask
your own question, for FREE!
Mathematics
14 Online
What is the probability of drawing a blue card, replacing it, and then drawing a blue card? http://i48.tinypic.com/35it1r8.jpg ~ Urgent HELPPPPP!
Still Need Help?
Join the QuestionCove community and study together with friends!
first draw: prob. = 3/5 2nd draw = 3/5 as the 2 draws are independent the probabilities are multiplied:" you required probability is 3/5 * 3/5 = 9/25
Oh my god! Thank you! I get it now! :)
good
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
12 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal